HUBBLE SPACE TELESCOPE DAILY REPORT #1336 PERIOD COVERED: 0000Z (UTC) 02/27/95 - 0000Z (UTC) 02/28/95 Daily Status Report as of 059/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of FOS/RD 5531 (FOS Target Acquisition Test) The Faint Object Spectrograph (Red Side) was used to make observations of standard star PG1115+080 to conduct acquisition tests. A TALED observation was also taken. HSTAR #5137 reported that there was a binary acquisition failure on FOS observation #14 at 058/0628Z. Observations #15 and 16 were also affected. No other anomalies were noted. 1.2 Completed Three Sets of WF/PC-2 5369 (HST Medium-Deep Survey: Cycle 4 Medium, High Galactic Latitude) The WF/PC-2 made a series of observations of high latitude objects as part of a medium deep sky survey. The observations were taken in parallel with other prime science observations. The observations were completed as scheduled, and no anomalies were noted. 1.3 Completed Two Sets of WF/PC-2 6140 (WF/PC-2 Cycle 4 Calibration: Ramp filters, Internal Calibration) The WF/PC-2 was used to make Earth calibration images as part of a ramp filter calibration test. The observations were completed as planned and no problems were reported. 1.4 Completed Two Sets of WF/PC-2 5479 (Subarcsecond Structure in Nearby AGNs) The WF/PC-2 was used to make snapshot observations of the nearby active galactic nuclei of MARK403 and MARK1239 to study the interactions between the central engine, the stars, and the interstellar matter in these AGNs. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed WF/PC-2 5490 (Exposing the Evolution of Distant Field Galaxies: Imaging in Mid UV) The WF/PC-2 was used to make observations of distant field galaxy 10HR-B. The objective is to study the morphology, interactions, and the star formation regions. The observations were executed as scheduled, and no anomalies were reported. 1.6 Completed FOC/96 5144 (Study of Optical Emission Associated With the Radio Jets and Hot Spots) The Faint Object Camera (f/96 ) was used to make observations of optical emissions from the jet and hot spots of galaxies NGC-1068. The observations were completed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 2 Successful : 2 Scheduled Reacquisitions : 12 Successful : 12 2.2 FHST Updates: Scheduled : 5 Successful : 5 2.3 Significant HSTARs: None 2.4 Operations Notes: Aside from the regularly-occurring data inversions while utilizing the STGT, there were no significant operational problems to report. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-